# EMFusion: Conditional Diffusion Framework for Trustworthy Frequency Selective EMF Forecasting in Wireless Networks

### v1

@@ V-C Evaluation Settings
For each day, the EMF time series contains $192$ regularly sampled observations. We use a 7-day lookback window, randomly sampled from the dataset, to predict the subsequent day. Formally, the historical input window is $H=7\times 192=1344$ time steps and the prediction window is $F=192$ steps. During sampling,we generate $\hat{N}=100$ scenarios for probabilistic forecasting. The hyperparameters for the EMFusion are set as follows: For training, the learning rate is 5e45e-4, the batch size is 64, and the model is trained for $1500$ epochs. The U-Net model has a depth of 6, and the attention mechanism uses $8$ heads. In the simulations, physical constants include the free-space wave impedance $Z_{0}=376.73\ \Omega$ and the antenna-gain calibration constant $A_{GC}=9.73$.
@@ V-D1 Comparative Analysis of EMFusion without External Conditions
TABLE V: Performance comparison Lower is better for CRPS, ND, NRMSE, MAE, MAPE, and Energy Score; for PICP, a value closest to 80% is preferred. Best per column in bold. Reported as mean $\pm$ standard error.
Method CRPS ND NRMSE MAE MAPE PICP Energy Score
NF 0.0110 $\pm$ 0.0070 0.1739 $\pm$ 0.0332 0.2447 $\pm$ 0.0941 0.0337 $\pm$ 0.0219 18.1875 $\pm$ 2.8910 69.1871 $\pm$ 6.0702 0.0163 $\pm$ 0.0103
IQLSTM 0.0120 $\pm$ 0.0076 0.1829 $\pm$ 0.0338 0.2575 $\pm$ 0.0954 0.0351 $\pm$ 0.0226 19.2144 $\pm$ 2.6370 61.8676 $\pm$ 5.9028 0.0192 $\pm$ 0.0120
WGAN 0.0123 $\pm$ 0.0071 0.1959 $\pm$ 0.0507 0.2588 $\pm$ 0.0998 0.0356 $\pm$ 0.0210 21.3099 $\pm$ 4.3154 50.0930 $\pm$ 9.8127 0.0202 $\pm$ 0.0114
VAE 0.0128 $\pm$ 0.0083 0.1908 $\pm$ 0.0411 0.2597 $\pm$ 0.0971 0.0367 $\pm$ 0.0237 20.6526 $\pm$ 3.8411 52.2893 $\pm$ 4.4282 0.0213 $\pm$ 0.0139
DDPM 0.0132 $\pm$ 0.0085 0.1989 $\pm$ 0.0520 0.2647 $\pm$ 0.0985 0.0385 $\pm$ 0.0260 22.5219 $\pm$ 6.0977 85.9521 $\pm$ 6.4352 0.0156 $\pm$ 0.0108
Dropout 0.0179 $\pm$ 0.0117 0.2418 $\pm$ 0.0481 0.2937 $\pm$ 0.0867 0.0470 $\pm$ 0.0299 30.2881 $\pm$ 6.8275 27.9310 $\pm$ 6.7606 0.0328 $\pm$ 0.0218
EMFusion-Uncond. 0.0212 $\pm$ 0.0154 0.2486 $\pm$ 0.0688 0.3257 $\pm$ 0.1048 0.0523 $\pm$ 0.0377 23.6121 $\pm$ 6.2094 33.0397 $\pm$ 10.4381 0.0389 $\pm$ 0.0284
EMFusion-MV(Multi) 0.0154 $\pm$ 0.0102 0.2083 $\pm$ 0.0412 0.2808 $\pm$ 0.0955 0.0411 $\pm$ 0.0266 22.8826 $\pm$ 5.1239 41.6521 $\pm$ 6.5349 0.0275 $\pm$ 0.0185
EMFusion-MV(Working Day) 0.0696 $\pm$ 0.0634 0.6345 $\pm$ 0.2315 0.6705 $\pm$ 0.2088 0.1433 $\pm$ 0.1262 62.0675 $\pm$ 22.9755 4.8629 $\pm$ 7.7988 0.1381 $\pm$ 0.1270
EMFusion-MV(Season) 0.0700 $\pm$ 0.0635 0.6355 $\pm$ 0.2347 0.6735 $\pm$ 0.2123 0.1445 $\pm$ 0.1265 61.9106 $\pm$ 23.5813 5.7850 $\pm$ 8.1614 0.1388 $\pm$ 0.1272
EMFusion-MV(WorkingHour) 0.0071 $\pm$ 0.0043 0.1049 $\pm$ 0.0241 0.1858 $\pm$ 0.0936 0.0195 $\pm$ 0.0120 10.7561 $\pm$ 1.4898 56.4799 $\pm$ 5.0838 0.0118 $\pm$ 0.0072
@@ VI Conclusion
Existing EMF forecasting approaches do not fully exploit the spatio-temporal dependencies across multiple frequency bands, operators, and contextual conditions. This study introduced EMFusion, a conditional diffusion based probabilistic forecasting framework. The framework employs conditional embeddings, cross-attention based context integration, and imputation-based sampling to improve temporal coherence. Experiments on real EMF datasets showed that EMFusion achieved lower deterministic forecasting errors and better probabilistic calibration than representative deep learning time-series and generative baselines across different operators and frequency selections, and that the predicted point values and prediction intervals were closer to the observed EMF trajectories. Future work may incorporate base-station location, transmit power, user distribution, and reconfigurable intelligent surface deployments as additional conditioning variables to further tighten the forecast distribution.
@@ Acknowledgment
The authors would like to thank Daniele Franci and Settimio Pavoncello for providing the EMF measurement hardware used to collect the Italian dataset.

### v2

@@ V-C Evaluation Settings
In our implementation, each 24-hour period of the EMF time series consists of $192$ regularly sampled observations. The model utilizes a 7-day lookback window to forecast the EMF trajectory for the subsequent day. Formally, this configuration yields a historical input window of $H=1344$ time steps and a forecast horizon of $F=192$ steps. The dataset is organized into sequential chunks, where each chunk comprises a 7-day lookback window ($H=1344$) and a 1-day forecast horizon ($F=192$). We allocate 60% of these chunks for training, with 20% reserved for validation and testing, respectively. The PICP confidence level is $\gamma=80\%$.
To construct the probabilistic prediction intervals during the inference phase, we generate an ensemble of $\hat{N}=100$ independent scenarios. The structural hyperparameters for EMFusion are optimized to capture complex temporal dependencies. Specifically, the diffusion process is configured with $T=200$ steps by default, while the U-Net backbone employs a hierarchical depth across its layers, with the dimension expansion factors $[1,2,2,2,2,2,2,2]$ respectively. For training, we employ a learning rate of $5\times 10^{-4}$ and a batch size of 64 over $1500$ epochs. The architecture further incorporates an 8-head attention mechanism within a U-Net model of depth 6 to ensure robust feature extraction. Finally, the simulation incorporates essential physical constants, including the free-space wave impedance $Z_{0}=376.73\ \Omega$ and the antenna-gain calibration constant $A_{GC}=9.73$. The training procedure incorporates standard regularization (e.g., early stopping based on validation loss, $L_{2}$ regularization, and batch normalization).
@@ V-D1 Comparative Analysis of EMFusion without External Conditions
TABLE V: Performance comparison Lower is better for ND, NRMSE, and MAPE; for PICP, a value closest to 80% is preferred. Best per column in bold. Reported as mean $\pm$ standard error.
Method ND NRMSE MAPE PICP
NF 0.1739 $\pm$ 0.0332 0.2447 $\pm$ 0.0941 18.1875 $\pm$ 2.8910 69.1871 $\pm$ 6.0702
IQLSTM 0.1829 $\pm$ 0.0338 0.2575 $\pm$ 0.0954 19.2144 $\pm$ 2.6370 61.8676 $\pm$ 5.9028
WGAN 0.1959 $\pm$ 0.0507 0.2588 $\pm$ 0.0998 21.3099 $\pm$ 4.3154 50.0930 $\pm$ 9.8127
VAE 0.1908 $\pm$ 0.0411 0.2597 $\pm$ 0.0971 20.6526 $\pm$ 3.8411 52.2893 $\pm$ 4.4282
DDPM 0.1989 $\pm$ 0.0520 0.2647 $\pm$ 0.0985 22.5219 $\pm$ 6.0977 85.9521 $\pm$ 6.4352
Dropout 0.2418 $\pm$ 0.0481 0.2937 $\pm$ 0.0867 30.2881 $\pm$ 6.8275 27.9310 $\pm$ 6.7606
EMFusion-MV Uncondition 0.2486 $\pm$ 0.0688 0.3257 $\pm$ 0.1048 23.6121 $\pm$ 6.2094 33.0397 $\pm$ 10.4381
Multi 0.2083 $\pm$ 0.0412 0.2808 $\pm$ 0.0955 22.8826 $\pm$ 5.1239 41.6521 $\pm$ 6.5349
WorkingDay 0.6345 $\pm$ 0.2315 0.6705 $\pm$ 0.2088 62.0675 $\pm$ 22.9755 4.8629 $\pm$ 7.7988
Season 0.6355 $\pm$ 0.2347 0.6735 $\pm$ 0.2123 61.9106 $\pm$ 23.5813 5.7850 $\pm$ 8.1614
WorkingHour 0.1049 $\pm$ 0.0241 0.1858 $\pm$ 0.0936 10.7561 $\pm$ 1.4898 56.4799 $\pm$ 5.0838
@@ V-D2 Univariate vs Multivariate Conditional EMFusion
Table IV shows a clear dominance of WorkingHour conditioning, while multi-conditioning often yields diminishing or negative returns. This reflects the characteristics of the University Hospital site, where EMF exposure closely follows visitor activity with a regular diurnal pattern, strong daytime peaks and low nighttime levels. As a result, WorkingHour captures the dominant temporal structure of BS power allocation and explains most of the variance in this dataset. Contextual relevance, however, is environment-dependent. In residential areas, traffic peaks may shift to evenings and weekends, increasing the importance of WorkingDay or holiday-related features. The performance degradation observed under multi-conditioning suggests that expanding the conditioning space without sufficient diversity in the training data can reduce learning efficiency. Higher-dimensional conditioning requires adequate coverage of contextual state combinations; when certain states (e.g., WorkingDay) are underrepresented in the dataset, the cross-attention mechanism struggles to prioritize the most informative signals.
@@
V-D3 Inference Complexity Analysis
Given a look-back window of length $H$, a forecast horizon $F$, an embedding dimension $d$, and $T$ total diffusion steps, we analyze the inference complexity as shown below. The inference complexity is dominated by the $T$ iterative denoising steps within the U-Net backbone and the cross-attention mechanism. For each denoising step, the complexity of the convolutional layers and attention projections scales with the total sequence length $(H+F)$ and the number of variates $N$. Specifically, the complexity of a single denoising step in our multivariate architecture is $\mathcal{O}(N\cdot(H+F)\cdot d^{2})$. Consequently, the total inference complexity for generating a multivariate forecast scales as $\mathcal{T}_{\text{inference}}=\mathcal{O}(T\cdot N\cdot(H+F)\cdot d^{2}).$ The inference time for EMFusion is in the order of milliseconds.
V-D4 Ablation Studies
TABLE VI: Ablation study on diffusion step size and U-Net depth in EMFusion. The training times reported in the tables are formatted as $H$:$M$:$S$, where $H$, $M$ and $S$ represents the number of hours, minutes, seconds, respectively.
Setting ND NRMSE MAPE PICP Time
Diffusion Steps 50 0.2393 0.3963 31.62 56.94 01:32:20
200 0.1335 0.2660 15.44 65.29 02:02:41
300 0.1759 0.3310 22.81 68.08 02:10:18
U-Net Depth 3 0.1786 0.3343 22.53 73.98 02:27:50
7 0.1872 0.3393 25.92 72.03 02:51:08
9 0.1625 0.3246 19.03 62.70 03:01:51
TABLE VII: Rolling window evaluation for EMFusion.
Period ND NRMSE MAPE PICP
2023/07/26 – 2023/12/12 0.1510 0.2040 20.25 55.17
2023/08/25 – 2024/01/11 0.1741 0.2374 18.16 56.69
2023/09/24 – 2024/02/10 0.1390 0.1954 15.84 59.76
2023/10/24 – 2024/03/12 0.1959 0.3737 21.96 61.17
Average 0.1650 0.2526 19.05 58.20
Full Dataset (No rolling window) 0.1335 0.2660 15.44 65.29
In Table VI, we varied the diffusion step size $\{50,200,300\}$ while fixing the U-Net depth at 8. Also, we investigated the impact of the U-Net backbone depth across $\{3,7,9\}$ layers while maintaining a fixed diffusion step size of $50$. Overall, performance variations across 50–300 steps are relatively small, indicating that the model is not highly sensitive to the diffusion steps. However, the depth of UNet architecture does impact the performance, i.e., the model with 9 layers outperform. This is because the deeper the model, the larger its receptive field across the entire time series, enabling it to capture abstract information within long sequences at the expense of additional complexity.
To evaluate the framework’s stability across distinct temporal segments, we then performed a rolling-window evaluation by sliding 5-month window across the 8-month period. The results in Table VII demonstrate consistent performance, confirming that EMFusion remains robust to the temporal variations. We note that the reduction in training data length inherent in the rolling-window setup can slightly impact performance; for instance, the average MAPE increases from 15.44% to 19.05% compared to the full dataset. This marginal degradation is attributed to the limited training data available within a window.
Figure 7: Training and validation loss across 2000 epochs for the W3 dataset.
To address the risk of potential overfitting, Fig. 7 illustrates the training and validation loss trajectories. The synchronized downward trend and stable convergence of both metrics demonstrate consistent generalization to unseen data without significant divergence. The model also includes explicit regularization through $L_{2}$ weight decay and Group Normalization, which help control model capacity and stabilize training. In addition, the diffusion objective perturbs inputs with Gaussian noise at multiple levels while sharing parameters across diffusion steps, introducing an implicit regularization effect that encourages robustness across noise conditions.
V-E Ethical and Deployment Considerations
V-E1 Data Governance
Reliable EMF forecasting requires robust data governance. In this study, data acquisition and analysis were conducted by the same organization, and the raw datasets were released as open source to ensure transparency and reproducibility. In multi-agency settings, strict compliance with regulations such as the General Data Protection Regulation (GDPR) [78] is essential, including the obfuscation of sensitive metadata and secure data-sharing protocols to prevent unauthorized access or misuse.
V-E2 Mitigating Monitoring Biases
Long-term monitoring can suffer from sampling biases due to sensor outages or irregular measurement intervals. EMFusion addresses this through its imputation-based sampling strategy (Section III-D), treating missing data as a structural inpainting task to preserve temporal coherence. While this reduces algorithmic bias, future monitoring systems should incorporate hardware-level detectors to identify abnormal sensor behavior or drift and maintain data reliability.
V-E3 Scalability for Nationwide Deployment
Nationwide deployment requires balancing computational overhead with monitoring density. The inference complexity of EMFusion scales linearly with the forecast horizon, $\mathcal{O}(F)$, enabling proactive management in large-scale deployments. However, national implementation also requires infrastructure for real-time data ingestion and long-term cloud storage.
@@ VI Conclusion
This paper introduced EMFusion, a conditional diffusion based probabilistic forecasting framework. The framework employs conditional embeddings, cross-attention based context integration, and imputation-based sampling to improve temporal coherence. Experiments have been conducted on real-world EMF dataset which showed that EMFusion achieved lower deterministic forecasting errors and better probabilistic calibration than representative deep learning time-series and generative baselines. EMFusion is a general data-driven forecasting framework that can be adapted to other locations if monitoring datasets and contextual information are available. Integrating principled data governance and scalable computational architectures will be important for transitioning EMFusion from a research framework to a reliable tool for nationwide regulatory compliance and public health monitoring. Future extensions may incorporate BS location, transmit power, and user distribution as additional conditioning variables to further tighten forecast distributions. Meta-heuristic optimization methods can also be applied for hyperparameter tuning to obtain more robust EMF forecasts [79].
@@ Acknowledgment
The authors would like to thank Dr. Daniele Franci and Settimio Pavoncello for providing the EMF measurement hardware used to collect the Italian dataset.